OpenStudy (anonymous):

factor using the difference of 2 squares 1.x^2-36 2.x^2+4 3.4x^2-25 4.y^2-1

OpenStudy (anonymous):

1.(x-6)(x+6)

OpenStudy (anonymous):

\[x^2+4x+4-4x=(x+2)^2-(2\sqrt{x})^2=(x+2-2\sqrt{x})(x+2+2\sqrt{x})\]

OpenStudy (anonymous):

3.(2x-5)(2x+5)

OpenStudy (anonymous):

4.(y_1)(y+1)

OpenStudy (anonymous):

ok so you guys are confusing me uzma its the same thing as before

OpenStudy (anonymous):

which 1?

OpenStudy (anonymous):

its gunna start out as factoring out the GCF and i have to do the difference of 2 squares method, or trial and error and i dont know how to do it.